A rectangle has an area of 1 over 6 and a length of 1.5 centimeters. What is the width and what is the perimeter? FRACTIONS
Answer by rfer(16322)

(1/6)/1.5=
(1/6)/(15/10)=
1/6*10/15=10/90=1/9 width
------------------------
2*1/6+2*1/9=
1/3+2/9=
3/9+5/9=8/9 perimeter